Suppose the velocity of money (V) is constant, Money supply (M) is growing at 5% per year, output (Y) is growing at 2% per year and interest rate (r) is 4%. Using the knowledge of the quantity theory of money and the Fischer effect;


a) Solve for i (nominal interest rate)


b) If the Central Bank increases the money supply by 2 percentage point per year, find the change in nominal interest rate


c) Suppose the growth rate of Y falls to 1% per year, what will happen to inflation? What can the Central Bank do if it wishes to keep inflation constant?



1
Expert's answer
2022-03-07T11:05:03-0500

Suppose V is constant, M is growing 5% per year, Y is growing 2% per year, and r = 4.


a)

We first find "\pi = 5 -2"


"\pi=3"


Therefore, "i = r + \pi"


"= 4 + 3"


"= 7"

b)

"\Delta i = 2"


It is just the same as the increase in the money growth rate.

"\u03c0=5\u22122"

"\u03c0=3"

Hence, "i = r + \pi"


"=4+3"


"= 7"


c)

The central bank does nothing "\Delta \pi =1". For the central bank to prevent inflation from rising, it must reduce the money growth rate by "1\%" point per year.